Grade 12th passIIT JEE Entrance Exam

S={1,2,3,.....,9} and A is a subset of S where n(A)=3 then sum of smallest element of all such possible sets 'A' is P then sum of digits of P is

kkji
10 Years agoGrade 12th pass

1 Answer

Vijay Mukati
10 Years ago
No. of Sets A, having 1 as smallest no. = 8C2, Addition = 1*8C2
No. of Sets A, having 2 as smallest no. = 7C2, Addition = 2*7C2
No. of Sets A, having 3 as smallest no. = 6C2, Addition = 3*6C2
No. of Sets A, having 4 as smallest no. = 5C2, Addition = 4*5C2
No. of Sets A, having 5 as smallest no. = 4C2, Addition = 5*4C2
No. of Sets A, having 6 as smallest no. = 3C2, Addition = 6*3C2
No. of Sets A, having 7 as smallest no. = 2C2, Addition = 7*2C2
Now add all this addition, you will get a number as P, Then find the sum of its digits.

Thanks.